In a ground state atom of S how many electrons total will have the quantum number ℓ = 0?


1
Expert's answer
2022-02-22T22:53:03-0500

l=2l=2


Magnetic quantum number from (-l \to +l)(−l→+l) :(m_1)=-2,-1,0,1,2:(m

1

)=−2,−1,0,1,2 (( Total =5)=5)


For magnetic quantum number,there are two spin quantum number that are (+\frac{1}{2},-\frac{1}{2})(+


So total number of electrons for\ l=2 \implies 5\times 2=10 l=2⟹5×2=10